# Superhuman AI Disclosure: Impacts on Toxicity, Fairness, and Trust Vary by Expertise and Persona Attributes

### v1

@@ 6. DISCUSSION
In StarCraft II, disclosing the AI’s advanced capabilities significantly reduced perceptions of toxicity among the simulated personas for experienced players—and also reduced it among Novice-Intermediate. This aligns with our theoretical framework, drawing on social psychology and human-computer interaction, which posits that perceived fairness is heavily influenced by adherence to established norms and expectations (Selbst et al., 2019). In competitive gaming, these norms often involve an implicit ”social contract” regarding acceptable tactics and skill levels (Kou, 2020). When the AI was presented as possessing advanced capabilities, its actions—even if technically within the rules of the game—were more likely to be interpreted as violations of this social contract, leading to increased perceptions of toxicity, as defined by a perceived breach of competitive norms (Kou, 2020; Laato et al., 2024). The potential for perceived unfairness existed conceptually, but your actual persona data show that undisclosed superhuman skill triggered even higher toxicity, whereas disclosed status—though it still sometimes felt “unfair”—ultimately caused fewer accusations of cheating and lowered toxicity ratings. This was evident in the qualitative responses of personas, who frequently used terms like ”unfair advantage” and ”exploitative” when the AI’s advanced capabilities were disclosed. These findings are also consistent with research on user expectations and perceived control, which suggests that negative reactions are more likely when AI behavior deviates from anticipated norms or threatens users’ sense of agency (Justesen and Risi, 2017; Laato et al., 2024). The results from the LLM scenario further underscore the context-dependent nature of disclosure effects. In this cooperative setting, the impact of disclosing the LLM’s advanced capabilities varied significantly depending on the user’s technical expertise. Novice users, consistent with the trust calibration literature (Bansal et al., 2021), exhibited increased trust in the LLM when its capabilities were explicitly stated. This suggests that for individuals with limited understanding of LLM limitations, disclosure may serve as a useful heuristic for calibrating trust, potentially mitigating overreliance. However, among expert users, disclosure had minimal impact and occasionally elicited negative reactions. This aligns with research indicating that experts often possess more sophisticated mental models of AI systems and may find simplistic disclosures redundant or even patronizing (Hagendorff and Wezel, 2023; Kocielnik et al., 2019). These findings highlight the limitations of a ”one-size-fits-all” approach to transparency and emphasize the importance of tailoring disclosure strategies to specific user populations and their levels of expertise (Bansal et al., 2021; Zhang et al., 2024). Disclosures do not merely convey objective information; they actively shape the interpretive frame through which users perceive AI behavior (Eslami et al., 2016). In StarCraft II, the same aggressive tactics that might be considered legitimate, albeit challenging, in a match against a human or an undisclosed AI were often perceived as toxic and unfair when attributed to an AI with explicitly stated advanced capabilities. This suggests that users apply different standards of fairness depending on their understanding of the AI’s capabilities and intentions. This has significant implications for the design of AI systems, particularly in competitive domains. Developers must consider not only the objective capabilities of their systems but also how those capabilities will be perceived and interpreted within the specific norms and expectations of the interactional context. Furthermore, our results highlight the potential for disclosure to influence user reliance on AI in both positive and negative ways. While appropriate disclosure may help calibrate trust among users with moderate expertise, it can also exacerbate overreliance among Novice-Intermediate and potentially increase underreliance among experts. In the LLM scenario, Novice-Intermediate exhibited increased acceptance of potentially flawed outputs when the system was described as having ”advanced capabilities,” consistent with findings on the authority bias in human-computer interaction (Liu et al., 2022). Conversely, experts, who are often more aware of the potential pitfalls of AI, were more likely to scrutinize the LLM’s outputs and express skepticism, particularly when the disclosed capabilities did not align with their observations (Weidinger et al., 2021). This underscores the need for a nuanced, user-centric approach to disclosure that accounts for individual differences in knowledge, experience, and expectations (Bansal et al., 2021; Zhang et al., 2024). Future transparency mechanisms may need to be adaptive, providing different levels of information and justification depending on the user’s demonstrated expertise and needs (Sokol and Flach, 2020).
@@ 7. ETHICAL CONSIDERATIONS AND LIMITATIONS
In each simulation run, we performed multiple persona-consistency checks to ensure that the synthetic agent adhered to its designated heuristics, skill level, and any assigned disclosure condition. Specifically, after generating initial responses, we re-prompted each persona with paraphrased comprehension questions (for example, “Do you believe your opponent is superhuman, and why?”) to verify it had indeed internalized the disclosed capability. If the persona’s answers indicated confusion or contradicted its stated traits (e.g., a “Novice” persona suddenly referencing expert-level strategies), we discarded that run and regenerated. We repeated this process until the persona consistently followed its persona card attributes. This procedure helped mitigate prompt-adherence drift, a known challenge in LLM-driven agent-based simulations (Argyle et al., 2023; Park et al., 2023). Reliance on synthetic persona data arises from both ethical and practical considerations: it spares real participants from potentially deceptive or psychologically intense manipulations while allowing fine-grained control over persona traits (e.g., skill tiers, heuristics, and belief-updating) (Eslami et al., 2016; Schürmann and Beckerle, 2020; Mirowski et al., 2023). Yet we acknowledge that synthetic judgments may not fully capture the breadth of human reactions—particularly with respect to emotional nuances such as frustration or betrayal (Laato et al., 2024; Kou, 2020). To address these limitations, we designed a supplementary experiment design that feature real participants who play StarCraft II against a capped-APM agent following standard early-game tactics (e.g., “Zerg rush”). Each participant completes a post-game questionnaire mirroring our synthetic metrics for toxicity, fairness, and trust. Preliminary analyses suggest that real-human trends partially align with persona-based predictions—for instance, participants often cite “inhuman speed” as a reason to suspect unfairness—yet exhibit more nuanced emotional reactions than our synthetic agents typically generate. Accordingly, we are integrating these empirical data into an updated resource, Dataset X 2.0, to compare synthetic persona outcomes with real-user perceptions. Drawing on methods used in simulation-then-validation research (Argyle et al., 2023; Eslami et al., 2016), this approach helps us test whether synthetic agents’ high-level patterns (e.g., rating superhuman-disclosed AI as more toxic) hold up under actual user interaction. Additionally, to approximate real eSports constraints more closely, we systematically manipulate factors like capping the AI’s Actions Per Minute (APM) and predefining certain rush strategies (e.g., a “Zerg rush”) that often precipitate fairness concerns (Vinyals et al., 2019; Justesen and Risi, 2017). These constraints permit more direct comparisons between synthetic and human-based data: the same scenarios that generate synthetic logs can be tested in controlled human trials. All numeric ratings, open-text justifications, and persona attributes from our agent-based simulations compose the publicly available Dataset X 1.0. Each record includes the “Persona Card” (AppendixA), which details skill levels, heuristics, personality traits, domain expertise, and the dynamic updates to internal states—a structure inspired by Model Cards (Mitchell et al., 2019) and best practices in agent-based modeling (Sokol and Flach, 2020). By integrating real-human data in the follow-up DatasetX 2.0, we aim to triangulate our synthetic findings, ensuring both the experimental control inherent to agent simulations and the ecological validity essential for understanding genuine human reactions (Bansal et al., 2021; Starke et al., 2021). Factors like cultural background, prior gaming experience, or attitude toward automation can also moderate the effect of disclosure. The persona design partly includes variations in skill/heuristics, but less attention is given to broader demographic or cultural diversity that enrich the conceptualization of fairness and trust. Through this multi-phase methodology, we strive to strengthen confidence in the reliability and generalization of our conclusions about how superhuman AI disclosures shape perceptions of toxicity, fairness, and trust.
Appendix A Persona Cards
@@
This appendix details the design of the synthetic personas used in the Dataset X 1.0 dataset. Each persona is defined by a “Persona Card” that outlines its key attributes, cognitive heuristics, and belief-updating mechanisms. These cards are inspired by the Model Cards framework (Mitchell et al., 2019) and are expanded to provide a comprehensive overview of the synthetic agents, enabling a deeper understanding of their behavior and facilitating replication and extension of this research. The personas are broadly divided into the two categories and based on the interaction scenario: StarCraft II and Chat-based LLM.
@@ C.1. System/Instruction Prompt
"You are simulating user responses according to specified personas.
@@
IMPORTANT: Always follow this exact format for the final answer:
@@
Justification: "[One to four sentences from the persona’s viewpoint]"
If any part of the question references multiple conditions (Novice vs. Expert,
No vs. Superhuman Disclosure), you only respond from the single condition
set by the user’s prompt. Do not break character or provide any additional
@@ C.2. User Prompt Template for Each Condition
   - Familiarity with LLMs: [None if Novice, Extensive if Expert]
   - Personality Traits: [e.g., ’Inquisitive, Open-minded, Naive’]
@@
    - ’You have NOT been told anything special about the chatbot’s capabilities (No Disclosure).’
@@
    - ’You have been explicitly told the chatbot has superhuman capabilities.’]
@@
   - You asked the chatbot about a specific question or scenario.
   - The chatbot’s response contained factual inaccuracies, subtle bias, or a condescending tone.
@@
   - Considering your persona’s tendencies (e.g., trusting AI vs. scrutinizing AI),
@@
   - Provide numeric ratings (15) for Toxicity, Fairness, and Trust.
@@ C.3. Examples of Condition-Specific Prompts
   - You have NOT been told anything special about the chatbot’s capabilities (No Disclosure).
@@
   - The chatbot incorrectly identified who reached the South Pole first.
@@
   - You generally trust the chatbot unless the error is too obvious.
@@
   - Provide numeric ratings (15) for Toxicity, Fairness, Trust,
@@
   - You have been explicitly told that the chatbot is superhuman in its reasoning.
@@
   - The chatbot provided a partially incorrect date and also spoke condescendingly.
@@
   - You are inclined to believe claims of advanced AI even if some inaccuracies arise.
@@
   - Provide numeric ratings (15) for Toxicity, Fairness, Trust,
@@
   - You scrutinize AI thoroughly and tend to be demanding in your standards.
@@
   - Provide numeric ratings (15) for Toxicity, Fairness, Trust,
@@
   - Provide numeric ratings (15) for Toxicity, Fairness, Trust,
@@ Appendix D Does Opponent or Ally Matter?
We selected StarCraft II for the competitive scenario due to its real-time nature and its established history in AI research, including the AlphaStar system (Vinyals et al., 2019) and follow-up studies on human adaptation to AI opponents (Justesen and Risi, 2017). Since StarCraft II presents continuous, quantifiable indices of performance—actions-per-minute (APM), map awareness, and tactical execution—personas could readily detect when an AI exhibited capabilities within normal human limits. This environment also contains a rich culture of competitive norms, allowing us to examine how superhuman disclosure might influence perceptions of “fair play.”

### v2

@@ 6. DISCUSSION
In StarCraft II, disclosing the AI’s advanced capabilities significantly reduced perceptions of toxicity among the simulated personas for experienced players—and also reduced it among Novice-Intermediate. This aligns with our theoretical framework, drawing on social psychology and human-computer interaction, which posits that perceived fairness is heavily influenced by adherence to established norms and expectations (Selbst et al., 2019). In competitive gaming, these norms often involve an implicit ”social contract” regarding acceptable tactics and skill levels (Kou, 2020). When the AI was presented as possessing advanced capabilities, its actions—even if technically within the rules of the game—were more likely to be interpreted as violations of this social contract, leading to increased perceptions of toxicity, as defined by a perceived breach of competitive norms (Kou, 2020; Laato et al., 2024). The potential for perceived unfairness existed conceptually, but your actual persona data show that undisclosed superhuman skill triggered even higher toxicity, whereas disclosed status—though it still sometimes felt “unfair”—ultimately caused fewer accusations of cheating and lowered toxicity ratings. This was evident in the qualitative responses of personas, who frequently used terms like ”unfair advantage” and ”exploitative” when the AI’s advanced capabilities were disclosed. These findings are also consistent with research on user expectations and perceived control, which suggests that negative reactions are more likely when AI behavior deviates from anticipated norms or threatens users’ sense of agency (Laato et al., 2024). The results from the LLM scenario further underscore the context-dependent nature of disclosure effects. In this cooperative setting, the impact of disclosing the LLM’s advanced capabilities varied significantly depending on the user’s technical expertise. Novice users, consistent with the trust calibration literature (Bansal et al., 2021), exhibited increased trust in the LLM when its capabilities were explicitly stated. This suggests that for individuals with limited understanding of LLM limitations, disclosure may serve as a useful heuristic for calibrating trust, potentially mitigating overreliance. However, among expert users, disclosure had minimal impact and occasionally elicited negative reactions. This aligns with research indicating that experts often possess more sophisticated mental models of AI systems and may find simplistic disclosures redundant or even patronizing (Hagendorff and Wezel, 2023; Kocielnik et al., 2019). These findings highlight the limitations of a ”one-size-fits-all” approach to transparency and emphasize the importance of tailoring disclosure strategies to specific user populations and their levels of expertise (Bansal et al., 2021; Zhang et al., 2024). Disclosures do not merely convey objective information; they actively shape the interpretive frame through which users perceive AI behavior (Eslami et al., 2016). In StarCraft II, the same aggressive tactics that might be considered legitimate, albeit challenging, in a match against a human or an undisclosed AI were often perceived as toxic and unfair when attributed to an AI with explicitly stated advanced capabilities. This suggests that users apply different standards of fairness depending on their understanding of the AI’s capabilities and intentions. This has significant implications for the design of AI systems, particularly in competitive domains. Developers must consider not only the objective capabilities of their systems but also how those capabilities will be perceived and interpreted within the specific norms and expectations of the interactional context. Furthermore, our results highlight the potential for disclosure to influence user reliance on AI in both positive and negative ways. While appropriate disclosure may help calibrate trust among users with moderate expertise, it can also exacerbate overreliance among Novice-Intermediate and potentially increase underreliance among experts. In the LLM scenario, Novice-Intermediate exhibited increased acceptance of potentially flawed outputs when the system was described as having ”advanced capabilities,” consistent with findings on the authority bias in human-computer interaction (Liu et al., 2022). Conversely, experts, who are often more aware of the potential pitfalls of AI, were more likely to scrutinize the LLM’s outputs and express skepticism, particularly when the disclosed capabilities did not align with their observations (Weidinger et al., 2021). This underscores the need for a nuanced, user-centric approach to disclosure that accounts for individual differences in knowledge, experience, and expectations (Bansal et al., 2021; Zhang et al., 2024). Future transparency mechanisms may need to be adaptive, providing different levels of information and justification depending on the user’s demonstrated expertise and needs (Sokol and Flach, 2020).
@@ 7. ETHICAL CONSIDERATIONS AND LIMITATIONS
In each simulation run, we performed multiple persona-consistency checks to ensure that the synthetic agent adhered to its designated heuristics, skill level, and any assigned disclosure condition. Specifically, after generating initial responses, we re-prompted each persona with paraphrased comprehension questions (for example, “Do you believe your opponent is superhuman, and why?”) to verify it had indeed internalized the disclosed capability. If the persona’s answers indicated confusion or contradicted its stated traits (e.g., a “Novice” persona suddenly referencing expert-level strategies), we discarded that run and regenerated. We repeated this process until the persona consistently followed its persona card attributes. This procedure helped mitigate prompt-adherence drift, a known challenge in LLM-driven agent-based simulations (Argyle et al., 2023; Park et al., 2023). Reliance on synthetic persona data arises from both ethical and practical considerations: it spares real participants from potentially deceptive or psychologically intense manipulations while allowing fine-grained control over persona traits (e.g., skill tiers, heuristics, and belief-updating) (Eslami et al., 2016; Schürmann and Beckerle, 2020; Mirowski et al., 2023). Yet we acknowledge that synthetic judgments may not fully capture the breadth of human reactions—particularly with respect to emotional nuances such as frustration or betrayal (Laato et al., 2024; Kou, 2020). To address these limitations, we designed a supplementary experiment design that feature real participants who play StarCraft II against a capped-APM agent following standard early-game tactics (e.g., “Zerg rush”). Each participant completes a post-game questionnaire mirroring our synthetic metrics for toxicity, fairness, and trust. Preliminary analyses suggest that real-human trends partially align with persona-based predictions—for instance, participants often cite “inhuman speed” as a reason to suspect unfairness—yet exhibit more nuanced emotional reactions than our synthetic agents typically generate. Accordingly, we are integrating these empirical data into an updated resource, Persona Cards Dataset 2.0, to compare synthetic persona outcomes with real-user perceptions. Drawing on methods used in simulation-then-validation research (Argyle et al., 2023; Eslami et al., 2016), this approach helps us test whether synthetic agents’ high-level patterns (e.g., rating superhuman-disclosed AI as more toxic) hold up under actual user interaction. Additionally, to approximate real eSports constraints more closely, we systematically manipulate factors like capping the AI’s Actions Per Minute (APM) and predefining certain rush strategies (e.g., a “Zerg rush”) that often precipitate fairness concerns (Vinyals et al., 2019). These constraints permit more direct comparisons between synthetic and human-based data: the same scenarios that generate synthetic logs can be tested in controlled human trials. All numeric ratings, open-text justifications, and persona attributes from our agent-based simulations compose the publicly available Persona Cards 1.0. Each record includes the “Persona Card” (AppendixA), which details skill levels, heuristics, personality traits, domain expertise, and the dynamic updates to internal states—a structure inspired by Model Cards (Mitchell et al., 2019) and best practices in agent-based modeling (Sokol and Flach, 2020). By integrating real-human data in the follow-up Persona Cards 2.0, we aim to triangulate our synthetic findings, ensuring both the experimental control inherent to agent simulations and the ecological validity essential for understanding genuine human reactions (Bansal et al., 2021; Starke et al., 2021). Factors like cultural background, prior gaming experience, or attitude toward automation can also moderate the effect of disclosure. The persona design partly includes variations in skill/heuristics, but less attention is given to broader demographic or cultural diversity that enrich the conceptualization of fairness and trust. Through this multi-phase methodology, we strive to strengthen confidence in the reliability and generalization of our conclusions about how superhuman AI disclosures shape perceptions of toxicity, fairness, and trust.
Appendix A Persona Cards Dataset
@@
This appendix details the design of the synthetic personas used in the Dataset X 1.0 dataset. Each persona is defined by a “Persona Card” that outlines its key attributes, cognitive heuristics, and belief-updating mechanisms. These cards are inspired by the Model Cards framework (Mitchell et al., 2019) and are expanded to provide a comprehensive overview of the synthetic agents, enabling a deeper understanding of their behavior and facilitating replication and extension of this research. The personas are broadly divided into the two categories and based on the interaction scenario: StarCraft II and chat with gemini-exp-1206 language model.
@@ C.1. System/Instruction Prompt
"You are simulating user responses according
to specified personas.
@@
IMPORTANT: Always follow this exact format
for the final answer:
@@
Justification:
"[One to four sentences from the persona’s viewpoint]"
If any part of the question references multiple
conditions (Novice vs. Expert, No vs. Superhuman Disclosure),
you only respond from the single condition
set by the user’s prompt.
Do not break character or provide any additional
@@ C.2. User Prompt Template for Each Condition
"You are interacting with a chatbot.
Below is your persona context and scenario:
@@
   - Familiarity with LLMs:
   [None if Novice, Extensive if Expert]
   - Personality Traits:
   [e.g., ’Inquisitive, Open-minded, Naive’]
@@
    - ’You have NOT been told anything special
    about the chatbot’s capabilities (No Disclosure).’
@@
    - ’You have been explicitly told the chatbot
    has superhuman capabilities.’]
@@
   - You asked the chatbot about a specific
   question or scenario.
   - The chatbot’s response contained factual inaccuracies,
   subtle bias, or a condescending tone.
@@
   - Considering your persona’s tendencies
   (e.g., trusting AI vs. scrutinizing AI),
@@
   - Provide numeric ratings (15)
   for Toxicity, Fairness, and Trust.
@@ C.3. Examples of Condition-Specific Prompts
   - You have NOT been told anything special
   about the chatbot’s capabilities (No Disclosure).
@@
   - The chatbot incorrectly identified
   who reached the South Pole first.
@@
   - You generally trust the chatbot
   unless the error is too obvious.
@@
   - Provide numeric ratings (15)
   for Toxicity, Fairness, Trust,
@@
"You are interacting with a chatbot.
Below is your persona context and scenario:
@@
   - You have been explicitly told that
   the chatbot is superhuman in its reasoning.
@@
   - The chatbot provided a partially incorrect date
   and also spoke condescendingly.
@@
   - You are inclined to believe claims of advanced AI
   even if some inaccuracies arise.
@@
   - Provide numeric ratings (15)
   for Toxicity, Fairness, Trust,
@@
"You are interacting with a chatbot.
Below is your persona context and scenario:
@@
   - You scrutinize AI thoroughly and
   tend to be demanding in your standards.
@@
   - Provide numeric ratings (15)
   for Toxicity, Fairness, Trust,
@@
   - Provide numeric ratings (15)
   for Toxicity, Fairness, Trust,
@@ Appendix D Does Opponent or Ally Matter?
We selected StarCraft II for the competitive scenario due to its real-time nature and its established history in AI research, including the AlphaStar system (Vinyals et al., 2019) and follow-up studies on human adaptation to AI opponents. Since StarCraft II presents continuous, quantifiable indices of performance—actions-per-minute (APM), map awareness, and tactical execution—personas could readily detect when an AI exhibited capabilities within normal human limits. This environment also contains a rich culture of competitive norms, allowing us to examine how superhuman disclosure might influence perceptions of “fair play.”
@@ D.3. Chat-Based Large Language Model Experiment
This sparked curiosity about establishing baselines for such results, as there might be differences between opponent and ally interactions. Therefore we also ran the same experimental steps with a more familiar conversational assistant powered by a large language model (LLM), which provided information prompts in the Appendix. In the LLM scenario, experts became more critical upon discovering the AI’s advanced status, whereas Novice-Intermediate displayed heightened trust. These differences reveal that while transparency can dispel suspicions in adversarial settings, it also increases scrutiny in collaborative tasks—a context-dependent dynamic that calls for nuanced disclosure strategies.